Solve each system by substitution.
-5x - 2y = 7
7x + y = -8

Answer by ikleyn(52802)

.
-5x - 2y =  7   (1)
 7x +  y = -8   (2)


Express y from the equation (2): y = -7x - 8, and substitute it into the equation (1).
You will get one single equation for x:

-5x - 2(-7x - 8) = 7.

Simplify and solve it:

-5x + 14x + 16 = 7,

9x = 7 - 16 = -9,

x = -1.

Now substitute the found value x = -1 into equation (2) and find y:

-7 + y = -8   ----->   y = -8 + 7 = -1.